OpenStudy (anonymous):

how would one do this problem?: (,/11 - ,/7)^2? it should be read as parenthesis, the square root of eleven minus the square root of 7 parenthesis squared

OpenStudy (anonymous):

satellite please go back to my question for a sec

OpenStudy (unklerhaukus):

\[(√11-√7)^2\] \[=11-2√11√7+7\] \[=18-2√77\]

OpenStudy (unklerhaukus):

did you get that formula right satellite?

OpenStudy (anonymous):

heck no i was wrong!!

OpenStudy (anonymous):

unkle has it , ignore me. i messed up

OpenStudy (unklerhaukus):

\[(a-b)^2=(a-b)(a-b)=a^2-2ab+b^2\]

OpenStudy (anonymous):

how do u get the 2 X ,/77? i tried foil, but it wont give me the same result

OpenStudy (anonymous):

nvm i gotcha

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

so you can leave the solution as =18−2√77 or =18-2√7√11

OpenStudy (anonymous):

okay thanks.